Development of Version 2.0 of a particular accounting software product is being considered by Jose​ Noguera's technology firm in Baton Rouge. The activities necessary for the completion of this project are listed in the following​ table:

                                                                                                                                

Activity

Normal Time​(weeks)

Crash Time​(weeks)

Normal Cost

Total Cost with Crashing

Immediate​ Predecessor(s)

A

44

33

$ 2 comma 000$2,000

$ 2 comma 650$2,650

long dash—

B

22

11

$ 2 comma 200$2,200

$ 3 comma 000$3,000

long dash—

C

33

33

$ 500$500

$ 500$500

long dash—

D

88

44

$ 2 comma 400$2,400

$ 2 comma 720$2,720

A

E

66

33

$ 1 comma 000$1,000

$ 1 comma 450$1,450

B

F

33

22

$ 3 comma 200$3,200

$ 4 comma 700$4,700

C

G

44

22

$ 1 comma 400$1,400

$ 1 comma 900$1,900

​D, E

​a) Based on the given information regarding the activities for the​ project, the project length​ =

1616

weeks.​b) The total cost required for completing this project on normal time​ =

​$1270012700.

​c) For reducing the duration of the project by one​ week, the activity that should be crashed first is activity

.The cost of the project based on the first activity selected for crashing will increase by

​$nothing.

​d) The maximum weeks by which the project can be reduced by crashing​ =

weeks.Total cost of crashing the project to minimum​ (or maximum weeks​ possible) =

​$nothing.
Answer #1

a) Critical path is the longest path = ADG = 4 + 8 + 4 = 16 weeks

b) Total cost = 2000 + 2200 + 500 + 2400 + 1000 + 3200 + 1400 = $12700

c) D has lowest crashing cost, hence can be crashed by 1 week at cost of (2720 - 2400) / 4 = $80

d) Crashing A , D and G to maximum i.e. A by 1 week, D by 4 week and G by 2 weeks.

Total cost = Crash cost - Normal cost = 2650 - 2000 + 2720 - 2400 + 1900 - 1400 = $1470

The new critical path is BEG = 2 + 6 + 2 = 10 weeks , so we need to reduce it by 1 week. Reduce E by 1 week at cost of (1450-1000)/3 = 150

Hence, total cost = 1470 + 150 = $1620 for project duration of 9 weeks.